You are considering how to invest part of your retirement savings. You have decided to put $ 500,000 into three​ stocks:

66 %of the money in GoldFinger​ (currently $ 27​/share), 19 % of the money in Moosehead​ (currently $ 83 ​/share), and the remainder in Venture Associates​ (currently $ 1 ​/share). Suppose GoldFinger stock goes up to $ 38​/share, Moosehead stock drops to $ 52 ​/share, and Venture Associates stock rises to $ 18 per share.

a. What is the new value of the​ portfolio? (Round to the nearest​ dollar.)

b. What return did the portfolio​ earn?

c. If you​ don't buy or sell any shares after the price​ change, what are your new portfolio​ weights?
Answer #1

1.
=500000*66%/27*38+500000*19%/83*52+500000*(1-66%-19%)/1*18
=1873962.51673360

2.
=1873962.51673360/500000-1
=274.7925033467200%

3.
Goldfinger=500000*66%/27*38/1873962.51673360=0.24784084
Moosehead=500000*19%/83*52/1873962.51673360=0.03176055
Venture Associates=1-0.24784084-0.03176055=0.72039861